the width of a rectangle is 8 cm less than its length. if the length is x cm and the area 105 cm^2. form an equation in x and solve

Answer by jorel1380(3719)

If x is the length, then the width would be x-8. So:
x(x-8)=105
x²-8x-105=0
(x+7)(x-15)=0
x=15 or -7
The length is 15 cm.; the width is 7 cm. ☺☺☺☺